OpenStudy (anonymous):

i need help on calc 2 sequences: Prove whether the sequence is convergent or divergent and if it is convergent, find it's limit: an= cos n/n!

OpenStudy (anonymous):

\[-1\le\cos n\le1~~\implies~~-\frac{1}{n!}\le\frac{\cos n}{n!}\le\frac{1}{n!}\]